methanol has a normal boiling point of 64.6°C and a heat of vaporization DHvap of 35.2kj/mol. What is the vapor pressure of methanol al 12.0°C?



1
Expert's answer
2021-02-25T03:44:09-0500

T1 = 12⁰C = 273+12 = 285K

T2 = 64.6⁰C = 273+64.6 = 337.6K

P2 = 760 torr

ln("\frac{P_2}{P_1}") = "\frac{\Delta H_{vap}}{R}" ( "\frac{1}{T_1}" - "\frac{1}{T_2}" )


ln("\frac{760Torr}{P_1}") = "\frac{35200 J mol^-¹}{8.314 J K^-¹ mol^-¹}" ( "\frac{1}{285K}" - "\frac{1}{337.6K}" )


ln("\frac{760Torr}{P_1}") = 4233.82 × 0.000546


ln("\frac{760Torr}{P_1}") = 2.311


"\frac{760Torr}{P_1}" = e2.311 = 10.08


P1 = "\frac{760Torr}{10.08}" =75.39 Torr